OpenStudy (anonymous):

solve system of equation: x+y=-1 y-z=4 2x+y+5z=-38

OpenStudy (anonymous):

i need step by step help please?

jhonyy9 (jhonyy9):

x+y=-1 y-z=4 2x+y+5z=-38 multiplie second by -1 and add to first and second x+y=-1 -y+z=-4 --------- x+z=-5 -y+z=-4 2x+y+5z=-38 -------------- 2x+6z=-42 divide both sides by 2 x+3z=-21 x+z=-5 x+3z=-21 subtract second from first ------------- 0 -2z = -5 -(-21) -2z = -5+21 -2z = 16 z= 16/(-2) = -8 x-8=-5 x=-5+8 x=3 x+y=-1 3+y=-1 y=-1-3 y=-4 x=3 y=-4 z=-8